OpenStudy (imamermaid29):

WILL FAN & MEDAL!! The system of equations is inconsistent. What are the missing values? 5x + ____y = 2 ____x + 3y = 8

OpenStudy (imamermaid29):

@jim_thompson5910

OpenStudy (anonymous):

We would need more information...

OpenStudy (imamermaid29):

thats all they give me .

OpenStudy (anonymous):

Not for her age

OpenStudy (anonymous):

oh, then i dont know... ask your teacher

OpenStudy (cresciez):

Do you have an idea of what the answer may be? Tell me it if you have one :)

OpenStudy (imamermaid29):

no i'm really bad at math XD

OpenStudy (cresciez):

Are there answer choices or is this all you got? ^-^

OpenStudy (cresciez):

And no worries, I'm bad at math too xD

OpenStudy (imamermaid29):

this is all i got :(

OpenStudy (anonymous):

The system and information is inconsistent xD

OpenStudy (cresciez):

Not exactly Savior, every question needs to have a solution unless there's an answer choice that says "no solution" but we don't have that right now

OpenStudy (cresciez):

Let's take small steps. Mermaid, do you know how to do these equations at all? Like have you done examples before?

OpenStudy (imamermaid29):

no this is my first time thats why i'm so confused

OpenStudy (cresciez):

Okay got it~ I found a website with a solution that may help and it gives us the steps. For convenience I'll just put it here :P So basically: There are infinite solutions for missing coefficients that will give an inconsistent system. 5x + by = 2 ax + 3y = 8

OpenStudy (imamermaid29):

im sooo confused XD

OpenStudy (cresciez):

LOL okay haha then let's try to find another

OpenStudy (cresciez):

This is pretty much the only approach to the solution. The variables change different letters but that's probably not important

OpenStudy (cresciez):

I'm just gonna give you the solution, I'm not sure how to explain. Have you studied the lesson? Here's the answer~ 5/a = b/3 ≠ 2/8 ab = 15, a≠20, b≠3/4

OpenStudy (cresciez):

@ganeshie8 can you help explain this to us? D: